OpenStudy (jennychan12):

When a neutron is removed from a 43/20 Ca atom of mass 42.958766 u, the latter will be transformed into a 43/20 Ca atom of mass 41.958618 u. What minimum energy is required for this removal?

OpenStudy (jennychan12):

It should be 

OpenStudy (anonymous):

E=mc^2 so c=3*10^8m/s. Plug in the values u will get the answer.

OpenStudy (anonymous):

Take the difference between the total final state masses and the total initial state mass (remember the mass of a neutron is 1.008982 AMU). When you use AMU's for your mass units then you can use the mass/energy conversion factor of 931.16 MeV per 1AMU. If the difference is negative then you must supply energy to make the reaction go. If it is positive the initial state was unstable and it would decay spontaneously.

OpenStudy (jennychan12):

if i use E = mc^2, is the mass the mass defect that i use?

OpenStudy (anonymous):

@jennychan12 mass = 42.958766 u - 41.958618 u =1.000148u......hey i just googled your question and so i got this formulae that needs to be applied but you better check with others too....

OpenStudy (jennychan12):

the E = mc^2 formula?

OpenStudy (anonymous):

@jennychan12 Yes i mean its usage......

OpenStudy (jennychan12):

ohh cool thanks :)

OpenStudy (anonymous):

@jennychan12 welcome